OpenStudy (anonymous):

if x = a + b , y =aw + bw^2 and z = aw^2 + bw Find x^3 + y^3 + z^3 =?

OpenStudy (anonymous):

x^3 + y^3 + z^3 = 3xyz ..... is this correct

OpenStudy (anonymous):

where W = cube root of unity.

OpenStudy (anonymous):

just want to check\[1+w+w^2=0\]\[w^2=-1-w\]right yahoo?

OpenStudy (anonymous):

if \(w=1\) its trivial and \[x^3+y^3+z^3=3(a+b)^3\]

OpenStudy (anonymous):

The answer should be \[3(a^3 + b^3)\]

OpenStudy (anonymous):

so ur problem is for when \(w\neq1\)

OpenStudy (anonymous):

but....there is no condition in the question)

OpenStudy (anonymous):

\[y^3+z^3=w^3[(a+bw)^3+(aw+b)^3]=(a+bw)^3+(aw+b)^3\]can u simplify this?

OpenStudy (anonymous):

Lol....Do u knw Such a Formula: x^3 + y^3 + z^3 = 3xyz

OpenStudy (anonymous):

no its not true

OpenStudy (anonymous):

Why.....

OpenStudy (anonymous):

\[x^3+y^3+z^3-3xyz=(x+y+z)(x^2+y^2+z^2-xy-xz-yz)\]

OpenStudy (anonymous):

Yup...nw i got it...thxx

OpenStudy (anonymous):

yahoo this time i really want u to go into it and do that simplifiction plz

OpenStudy (anonymous):

yes....i am Doing it....

OpenStudy (anonymous):

\[a^3 + 3a^2bw + 3ab^2w^2 + b^3 w^3 + a^3w^3 + 3a^2w^2 + 3awb^2 +b^3\]

OpenStudy (anonymous):

\[w^3 = 1\]

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

ok replace \(w^3\) with 1 and \(w^2\) with \(-1-w\)

OpenStudy (anonymous):

simplify and tell me what u get

OpenStudy (anonymous):

\[2a^3 + 2b^3 +3ab^2w - 3ab^2 - 3a^2 - 3a^2w\]

OpenStudy (anonymous):

u most get this\[a^3 + 3a^2bw + 3ab^2w^2 + b^3 w^3 + a^3w^3 + 3a^2w^2b + 3awb^2 +b^3\\=2(a^3+b^3)-3a^2b-3ab^2\]and finally\[(a+b)^3+2(a^3+b^3)-3a^2b-3ab^2=3(a^3+b^3)\]

OpenStudy (anonymous):

yes.....i would have made a mistake.....thxx.. \[\boldsymbol{\mathfrak{ MUKUSHLA}}\]

OpenStudy (anonymous):

\[\large \color\green{\text{Welcome}}\]

OpenStudy (anonymous):

i syudied..latex today.....

OpenStudy (anonymous):

Which website...u use for it...link..plzz

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=simplify+%28a+%2B+b%29^3+%2B+%28b+E^%28%282+I+ \[Pi]%29%2F3%29+%2B+++++a+E^%28%284+I+\[Pi]%29%2F3%29%29^3+%2B+%28a+E^%28%282+I+\[Pi]%29%2F3%29+%2B+++++b+E^%28%284+I+\[Pi]%29%2F3%29%29^3

OpenStudy (anonymous):

\[ (a+b)^3+\left(e^{\frac{2 i \pi }{3}} a+e^{-\frac{1}{3} (2 i \pi )} b\right)^3+\left(e^{-\frac{1}{3} (2 i \pi )} a+e^{\frac{2 i \pi }{3}} b\right)^3=\\ 3 a^3+3 e^{\frac{2 i \pi }{3}} a^2 b+3 e^{-\frac{1}{3} (2 i \pi )} a^2 b+3 a^2 b+3 e^{\frac{2 i \pi }{3}} a b^2+3 e^{-\frac{1}{3} (2 i \pi )} a b^2+3 a b^2+3 b^3=\\ 3 a^3+3 e^{\frac{2 i \pi }{3}} a^2 b+3 e^{-\frac{1}{3} (2 i \pi )} a^2 b+3 a^2 b+3 e^{\frac{2 i \pi }{3}} a b^2+3 e^{-\frac{1}{3} (2 i \pi )} a b^2+3 a b^2+3 b^3=\\ 3 a^3+3 a^2 b+6 a^2 b \cos \left(\frac{2 \pi }{3}\right)+3 a b^2+6 a b^2 \cos \left(\frac{2 \pi }{3}\right)+3 b^3= \\ 3 a^3+\frac{6}{2} (-1) a^2 b+3 a^2 b+\frac{6}{2} (-1) a b^2+3 a b^2+3 b^3=3 a^3+3 b^3 \]